# Multi-Shot Character Consistency for Text-to-Video Generation

### v1

@@ Abstract
Text-to-video models have made significant strides in generating short video clips from textual descriptions. Yet, a significant challenge remains: generating several video shots of the same characters, preserving their identity without hurting video quality, dynamics, and responsiveness to text prompts. We present Video Storyboarding, a training-free method to enable pretrained text-to-video models to generate multiple shots with consistent characters, by sharing features between them. Our key insight is that self-attention query features (Q) encode both motion and identity. This creates a hard-to-avoid trade-off between preserving character identity and making videos dynamic, when features are shared. To address this issue, we introduce a novel query injection strategy that balances identity preservation and natural motion retention. This approach improves upon naive consistency techniques applied to videos, which often struggle to maintain this delicate equilibrium. Our experiments demonstrate significant improvements in character consistency across scenes while maintaining high-quality motion and text alignment. These results offer insights into critical stages of video generation and the interplay of structure and motion in video diffusion models. Project page: https://research.nvidia.com/labs/par/video_storyboarding/.
Figure 1: Consistent Video Storyboarding. (click-to-view-online) Our method generates video shots from input prompts, ensuring consistent subjects across shots.
@@ 1 Introduction
Generating videos from text prompts is advancing rapidly, but it is still not feasible to create long coherent video sequences. A natural alternative would be to generate multiple short videos that share the same characters. Indeed, cinematic videos typically consist of many shorter shots, making them more engaging. The challenge is that although current text-to-video (T2V) models excel at generating individual clips, they struggle to maintain character consistency across multiple scenes. We wish to generate multiple video shots of the same characters, preserving their identity across all video scenes, while preserving video quality and dynamics. Such technology would open new opportunities in content creation storytelling and entertainment.
Recently, several methods (Tewel et al., 2024; Fan et al., 2024) have been proposed to generate images with consistent characters across various text prompts. However, achieving consistency in video is inherently more challenging due to a fundamental conflict between maintaining character identity and ensuring dynamic motion. In video, the same features often encode both identity and motion. Thus, when motion occurs, models could interpret it as a change in identity. This problem is unique to video and has not been explored in the context of images. Consequently, existing image-based consistency methods struggle to generalize to video, failing to achieve both character consistency and dynamic motion simultaneously.
We present Video Storyboarding, a training-free method for generating multi-shot videos with consistent characters. Our approach uses pre-trained T2V models by sharing features between the video shots. We first demonstrate that the self-attention query (Q) components primarily encode motion information, but they also contain identity features of the generated characters. When features are shared, injecting inconsistent Q components across videos preserves motion but disrupts character identity. Conversely, maintaining similar Q components across generated videos ensure character identity but unifies motion.
To address this, we propose a two-phase approach (Fig. 2): Q-Preservation followed by Q-Flow. In the Q Preservation phase, we maintain motion structure by replacing our Q values with those from “vanilla” (unconstrained) video generation. Then, the subsequent Q-Flow phase aims to maintain the optical flow of vanilla queries rather than their exact values. It employs flow maps derived from vanilla key frames to guide the injection of our identity-preserving Q features. This ensures character identity is maintained by placing consistent features in motion-appropriate locations. Finally, we combine this with a frame selection strategy for extended attention, that promotes visual coherence without freezing motion.
Through extensive ablation studies and comparisons with baseline methods, we show that Video Storyboarding significantly improves character and object consistency across scenes while maintaining high-quality motion and adhering to the input text prompt. Our ablation studies provide insights into the critical stages of video generation, the relationship between structure and motion, and the impact of different consistency strategies on the final output.
Our contributions are: (1) Video Storyboarding, a novel training-free method that enables character consistency in generating multi-shot video sequences, while maintaining motion adherence to prompts. (2) We reveal the dual role of self-attention query (Q) features in encoding both  motion and identity information . (3) We propose a novel two-phase query injection mechanism to balance these aspects, addressing the unique challenges posed by the temporal dimension. (4) We demonstrate significant improvements in character consistency while maintaining motion quality over baselines, tested with two pretrained models.
@@ 2 Related work
In our work, we explore bringing the training-free, feature-sharing consistency approach to the realm of video generation, with the goal of maintaining a consistent character across video scenes.
Figure 2: Video Storyboarding Architecture: Our consistent denoising process has two phases: Q Preservation and Q Flow. We first generate and cache video shots using “vanilla” VideoCrafter2. In Q Preservation ($T\rightarrow t_{pres}$), we use Vanilla Q Injection to maintain motion structure by replacing our Q values with vanilla ones. In Q Flow ($t_{pres}\rightarrow t_{0}$), we use a flow map from vanilla key frames to guide Q feature injection. This phase maintains character identity by allowing the use of Q features from our consistent denoising process, while the flow map ensures that these identity-preserving features are applied in a way that’s consistent with the original motion. Throughout, we employ two complementary techniques: framewise subject-driven self-attention for visual coherence, and refinement feature injection (Section 4.3) to reinforce character consistency across diverse prompts.
@@
Attention-based consistency. When using text-to-image models to generate (Wu et al., 2023; Ceylan et al., 2023; Khachatryan et al., 2023) or modify a video (Geyer et al., 2023), an extended self-attention block (Wu et al., 2023) is often employed to share keys and values across different frames, enabling them to draw visual appearances from each other and enhance consistency. Beyond cross-frame consistency, it has been used to inject consistent identities from a source image to video (Xu et al., 2023; Hu et al., 2023; Chang et al., 2023; Tu et al., 2023), maintain appearance in layout editing (Cao et al., 2023; Avrahami et al., 2024), combine appearances (Alaluf et al., 2023), for personalization (Gal et al., 2024; Zeng et al., 2024) and style transfer (Hertz et al., 2023).
In Consistory, Tewel et al. (2024) demonstrate that masked extended-attention can promote subject consistency across batch-generated images. We extend this to video generation, maintaining character consistency between video clips. We also use a masked extended-attention mechanism, but couple it with query-injection components to better maintain motion.
3 Preliminaries
3.1 Notations

### v2

@@ Abstract
Text-to-video diffusion models have shown remarkable progress in generating coherent video clips from textual descriptions. However, the interplay between motion, structure, and identity representations in these models remains under-explored. Here, we investigate how self-attention query features (a.k.a. $Q$ features) simultaneously govern motion, structure, and identity and examine the challenges arising when these representations interact. Our analysis reveals that Q affects not only layout, but that during denoising Q also has a strong effect on subject identity, making it hard to transfer motion without the side-effect of transferring identity. Understanding this dual role enabled us to control query feature injection (Q injection) and demonstrate two applications: (1) a zero-shot motion transfer method that is 20$\times$ more efficient than existing approaches, and (2) a training-free technique for consistent multi-shot video generation, where characters maintain identity across multiple video shots while Q injection enhances motion fidelity.
Project page: https://research.nvidia.com/lab/par
/MotionByQueries/
Figure 1: (click-to-view-online) Our analysis reveals differences in Q-injection between text-to-video and text-to-image models. One key observation is that in text-to-video models, zero-shot Q injection can transfer structure and motion between different video shots. However, when a target video is prompted for the same subject, Q injection suffers from identity leakage.
@@ 1 Introduction
Video generation from text is at the forefront of generative AI. Despite great progress in controlling entities in video, several major challenges remain, including generating natural and engaging motion and preserving consistent identity of entities throughout the video. These two goals tend to form a trade-off: It is easy to preserve consistency if motion is strongly limited, and making entities move makes it harder to enforce consistency, because the appearance of an entity changes. A major challenge lies in understanding how motion and identity are represented in various components of video generation models and how to effectively control them.
This limited understanding hinders downstream applications in video generation. For instance, current motion transfer approaches Yatim et al. (2024); Zhao et al. (2024); Wang et al. (2024a) still rely on tuning or test-time optimization, while their text-to-image layout transfer counterparts already operate through inference-time feature manipulation Cao et al. (2023); Alaluf et al. (2023). Better model understanding could potentially lead to similar strides here. As another example, consider consistent characters in multi-shot video generation, where the goal is to preserve consistency of character identity and appearance across various shots. Image-based models tackle this task through feature-sharing, but applying the same ideas to video leads to loss of motion because the shared features encode both identity and motion.
To understand representation of motion and identity in a video model, let us first draw analogies from text-to-image (T2I) models which are better understood. Motion can be viewed as a ”3D-shape” in the 3D tensor defined by a sequence of frames, so it is natural to look at representation of shape and structure in image generation models.
Previous works showed that diffusion-based T2I models establish the layout of the generated image in early generation steps (Patashnik et al., 2023). Several studies Cao et al. (2023); Alaluf et al. (2023); Tewel et al. (2024) have analyzed how different model components determine this structure. They showed that self-attention queries (Q) encode structural layout information. They found that injecting queries between images during generation “copies” shape while preserving appearance. As motion is the video equivalent of structure, we investigate the relationship between query vectors, motion and identity in video generation.
We conduct an empirical analysis, and observe that in contrast to image models, Q in video-generation models affects both motion and identity. Moreover, videos require more denoising steps than images to capture motion patterns. We use this insight to control motion in two different applications: motion transfer and consistent multi-shot video generation.
For motion transfer, we find that injecting Q features from a source video during denoising of a generated video, allows to transfer the motion to a new video in a zero-shot manner (without any additional fine-tuning or optimization). Our simple pipeline achieves generation quality close to leading methods, while being 20$\times$ more efficient than existing approaches.
For consistent multi-shot video generation, we build on insights from multi-shot image generation Tewel et al. (2024) relying on shot-to-shot extended attention. We find that Q injection from unconstrained video generation can preserve layout and motion diversity. However, unlike images, video generation requires more Q injection steps, causing identity leakage from the unconstrained source video and compromising shot-to-shot consistency. We address this with a two-phase approach: Q-Preservation maintains motion structure using Q values from unconstrained generation, while Q-Flow instead preserves feature flow maps, avoiding leaked identity in later steps.
Our main contributions: 1) We provide the first systematic analysis of Q-features in text-to-video diffusion models, revealing their dual role in encoding both motion and identity information, with effects persisting longer into the denoising process. 2) We introduce “Motion by Queries”, a zero-shot motion transfer approach that achieves comparable quality to current methods while being 20$\times$ more efficient. 3) We present the first training-free method for consistent multi-shot video generation that balances character consistency and motion quality.
Figure 2: Same-class motion transfer suffers from identity leakage (purple), which worsens with increased Q injection. Cross-class transfer (green) achieves reasonable separation at 40% injection, where motion quality is also preserved. The leftmost purple data point corresponds to results for random images of the same-class, as no injection occurs. Quantitative results were obtained using the motion-transfer data of Wang et al. (2024a). For illustration we show frames of the videos in Fig. 1.
@@ 2 Related work
Motion Transfer in Video Generation. Motion transfer aims to guide video generation using motion patterns extracted from a source video, spanning camera movements and object deformations while maintaining structural integrity. Recent approaches commonly require model adaptation, either through fine-tuning temporal attention layers (Jeong et al., 2024; Ren et al., 2024), employing dual-path LoRA architectures to separate appearance and motion learning (Zhao et al., 2024), or learning specialized motion embeddings (Wang et al., 2024a). Alternative methods utilize test-time optimization to guide the generation process (Yatim et al., 2024). In contrast to these approaches that require optimization or fine-tuning steps, our method enables zero-shot motion transfer through query injection, requiring no additional training or optimization.
Video Editing. While motion transfer focuses on replicating specific motion patterns from source to target videos, video editing methods focus on modifying specific attributes of an input video while preserving its other characteristics. Recent approaches employ various mechanisms for content modification: Bai et al. (2024) is also based on feature injection. However, it focuses on very specific attribute changes while preserving appearance and lacks the broader context of our Q injection analysis. Zhao et al. (2023) combine ControlNet with CLIP-based guidance. Earlier works adapted text-to-image models for video editing through different attention manipulation strategies, such as injecting complete self-attention maps (Qi et al., 2023). While this direction spawned numerous works (Liu et al., 2024; Shin et al., 2024; Cong et al., 2023; Wu et al., 2023), these image-based approaches cannot faithfully capture motion patterns without replicating the source appearance, as the underlying models lack the capacity to process temporal dynamics independently.
@@
Structure and Appearance in Text-to-Image Models Early work in text-to-image generation revealed distinct roles for different components of attention mechanisms in diffusion-based models. Tewel et al. (2023) demonstrated that attention maps, derived from Keys and Queries, act as a “Where” pathway controlling the compositional structure, while Values serve as a “What” pathway determining visual appearance. Further studies established that the layout of generated images is primarily determined in the initial denoising steps (Patashnik et al., 2023). Through detailed analysis of model components,  (Alaluf et al., 2023; Tewel et al., 2024) identified self-attention queries as crucial encoders of structural layout information. Our work extends these insights to video generation, revealing that queries play a more complex role than previously understood - affecting both motion and identity. Additionally, we find that capturing motion patterns in videos requires more denoising steps compared to image generation, highlighting fundamental differences between these domains.
@@
Extended Attention Sharing. When using text-to-image models to generate (Wu et al., 2023; Ceylan et al., 2023; Khachatryan et al., 2023) or modify a video (Geyer et al., 2023), an extended self-attention block (Wu et al., 2023) is often employed to share keys and values across different frames, enabling them to draw visual appearances from each other and enhance consistency. Beyond cross-frame consistency, it has been used to inject consistent identities from a source image to video (Xu et al., 2023; Hu et al., 2023; Chang et al., 2023; Tu et al., 2023), maintain appearance in layout editing (Cao et al., 2023; Avrahami et al., 2024), combine appearances (Alaluf et al., 2023), for personalization (Gal et al., 2024; Zeng et al., 2024) and style transfer (Hertz et al., 2023).
3 Analysis: The role of Query features in text-to-video generation
To study the role of Q-features in text-to-video (T2V) models, we design an injection experiment. The idea is simple: we take a source video $V_{S}$ with a known text description $\tau_{S}$, and also generate a target video $V_{T}$, using a text prompt $\tau_{T}$. We then record Q features from $V_{S}$, inject them into the generation of $V_{T}$, and analyze the effect of that injection.
More specifically, Given the video $V_{S}$ we add noise to it with a level that correspond to some noisy step $t$, yielding a noisy latent $z_{t}$. We then perform a single DDPM denoising step (with a $50$ step schedule), and record the spatial Q features of all the self attention layers of the diffusion model. We repeat this $20$ times for various noise levels. At the end of this procedure, we have a sequence of $20$ $Q$ tensors ($Q_{S}(50),Q_{S}(49),...,Q_{S}(30)$ DDPM steps corresponding to t=1000, 980, 960, …, 600). Finally, we generate a new video $V_{T}$ with its prompt $\tau_{T}$, while injecting $Q_{S}(t)$ tensors at the first k DDPM denoising steps. We vary the amount of denoising steps that receive Q-injection: From no injection at all (0%) to injecting $Q$ through 40% of DDPM steps.
To understand the effect of Q-injection, we now measure similarity between source and target videos in two aspects: Identity Leakage and Motion Fidelity. Identity Leakage measures the mean DINO similarity between the frames of the source $V_{S}$ and target $V_{T}$ videos. Motion Fidelity Yatim et al. (2024) measures cross-correlation between point tracks in source and target videos:
$\frac{1}{m}\sum_{\tilde{p}\in\tilde{\mathcal{P}}}\underset{p\in\mathcal{P}}{\text{max}}\ \textbf{corr}(p,\tilde{p})+\frac{1}{n}\sum_{{p}\in{\mathcal{P}}}\underset{\tilde{p}\in\tilde{\mathcal{P}}}{\text{max}}\ \textbf{corr}(p,\tilde{p})$ (1)
where $\mathcal{P}=\{p_{1},\dots,p_{n}\},\tilde{\mathcal{P}}=\{\tilde{p}_{1},\dots,\tilde{p}_{m}\}$ are point tracks in source and target videos.
We present our analysis results in Fig. 2. We analyze two Q-injection setups: one where the source video prompt $\tau_{S}$ and target prompt $\tau_{T}$ share the same subject (purple) and another where they have different subjects (green). We discuss three key findings, highlighting the differences in how Q vectors behave in T2V models compared to T2I models.
First, Fig. 2 (bottom-left), illustrates that Motion Fidelity grows monotonically with Q injection duration and reaches high similarity at 40% injection. We find that, unlike images–where structure is established early in the denoising process–videos require significantly more steps to set the motion structure. Second, more surprising is the top-left panel. It reveals that identity similarity also grows monotonically with duration of injecting Q. This suggests that video generation models also encode identity information into the Q vectors–an intriguing shift from its traditionally assumed role in T2I models. Third, we observe an interesting phenomenon: when $\tau_{S}$ and $\tau_{T}$ use the same subject, the target video often features a subject with an appearance identical to that of the source video while maintaining the background specified in $\tau_{T}$. This identity “leakage” is significantly less pronounced when $\tau_{T}$ and $\tau_{S}$ feature different subjects. Qualitatively, in Fig. 2 (right), we present a source video of “A horse galloping in the savanna” with a distinct white spot on its back. Notice that as the number of injection steps increases, the target horse (purple) becomes more similar to the source horse while preserving the cloud background specified in the prompt. On the other hand, when $\tau_{T}$ describes a giraffe, its shape and motion become more similar to those of the source horse, but its appearance remains that of a giraffe. This suggests a clear distinction in Q-injection behavior: it transfers appearance and motion when the subjects match, while for different subjects, the transfer primarily affects motion characteristics.
Figure 3: Motion is compromised with extended attention across video-shots. To recover the motion, longer q-injection periods are required, which consequently increases identity leakage.
Finally, we examine the effect of Q-injection when the target consists of multiple videos sharing generation features. Feature sharing within a batch is a widely used approach in T2I generation, aiding tasks such as consistent character generation  (Tewel et al., 2024) and editing  (Cao et al., 2023). A common technique is the extended attention mechanism, which modifies the model’s self-attention layers to attend across elements of the multiple elements in the same batch. Here, we explore the impact of Q-injection when generating a batch of 3 videos at a time and utilizing the shared Extended Attention between the video shots. In Fig. 3, we include the Extended Attention curve to visualize motion fidelity in this scenario. Our results show that extended attention induces a ”motion freeze” effect, requiring Q to be injected for more steps to achieve the same motion as the source video. Moreover, since Q-injection in the Extended Attention setting requires more steps, it also introduces identity information, making it difficult to disentangle motion injection from identity injection.
4 Application 1: Motion transfer
Our first application is motion transfer. Motion transfer allows creators to control the motion in a fine way, which cannot be achieved using text. In this problem we are given a video $V_{S}$ that contain some pattern of movement, and we wish to generate a new video $V_{T}$ that follows the same movement patterns. The movement can be of entities in the video, or the camera or both.
Figure 4: (click-to-view-online) Qualitative Results, Motion Transfer. Each pair of rows are frames from source (pair-top) real video and target (pair-bottom) generated video. Transferring Motion by Queries allows to use source videos to inject camera motion (top), non-rigid movement (middle), and combinations of movements (bottom). See the supplemental for videos and more examples.
Figure 5: (click-to-view-online) Qualitative Comparisons, Motion Transfer.
4.1 Method for Motion transfer
Our approach, called “Motion by Queries” follows the experiment described in section 2. First, we extract a series of Q-features from $V_{S}$, by denoising the model to various timesteps ($t=1000,980,\ldots,600$), obtaining $[Q_{S}(50),\ldots,Q_{S}(30)]$. Then, we inject those Q features during the generation of the target video $V_{T}$, this time with the new prompt $\tau_{T}$.
Implementation details: We only record and inject the Q features of the conditional forward pass. To ensure consistency with the video model’s latent space, we follow Meiri et al. (2023) and add minor noise to the source video latent (two steps) and denoise it back to a new clean latent before extracting Q-features. For recording Q features, we use the same noise seed across different noise levels. During generation, we start from a random initialization. For our experiments, we use VideoCrafter2 (Chen et al., 2024).
4.2 Experiments
We evaluate Motion by Queries qualitatively and quantitatively using a Motion-Transfer benchmark Wang et al. (2024a), comprising 66 video-edit text pairs, guided by 22 source videos. Videos were sourced from DAVIS Perazzi et al. (2016), WebVID Bain et al. (2021), and online resources, representing diverse scenes, objects, and motion. For the base text-to-video model, we employ VideoCrafter2 Chen et al. (2024).
Baselines: We compare our approach with recent methods: (1) Diffusion Motion Transfer (DMT), a test-time optimization approach Yatim et al. (2024), and (2) Motion-Inversion (MI) Wang et al. (2024a), a fine-tuning-based concurrent method that learns motion-specific embeddings that modulate temporal attention activations from the source video’s motion. MI indicates results we reproduced from their online code. Except for one case in the supplemental, all qualitative comparisons use videos from MI and DMT project pages. We also adopt from Wang et al. (2024a) the quantitative results of Video Motion Customization (VMC) Jeong et al. (2024) and Motion-Director (MD) Zhao et al. (2024), both are fine-tuning-based approaches.
Evaluation Metrics: We measure Identity Leak (ID. Leak) from source to target video, as described in Section 3. Following Wang et al. (2024a), we also measure Motion Fidelity (MF), Text Similarity (Text), and Temporal Consistency (Temp. C.). MF measures point track correlation between generated and source videos, as described in Section 3. Text Similarity is the average CLIP-Text score between generated frames and the textual prompt Radford et al. (2021). Temporal Consistency Huang et al. (2024) measures video coherence by averaging CLIP-Image similarity between consecutive frames in the generated video.
Additionally, we compare the runtimes of different methods and their relative overhead compared to the base video model each method uses. We break down the runtime into its components: “Source Invers.” is the time for DDIM inversion in baselines or recording Q features in our approach. “Optim./Tuning” is the optimization or fine-tuning time, and “Infer” is the time to generate a target video. Note that in DMT, Infer and Optim are done jointly. “Sum” is the total time for each method, and “Overhead” is the ratio between total time and inference time in the base model. Finally, runtimes were evaluated using an NVIDIA H100 GPU, when generating a video of 24 frames at a resolution of 576$\times$320.
Quantitative Results: In terms of performance, our approach demonstrates competitive results across various metrics while maintaining simplicity and optimization-free advantages (Tables 1, 2). Specifically, our method achieves lower (better) identity leakage (38.6) compared to MI (43.7), while text alignment (28.8) and temporal consistency (97.0) perform on par with baselines. Regarding motion fidelity (MF), our method achieves 91.5. Although lower than MI (97.0), fidelity remains adequate for many practical applications, as supported by the qualitative results. Notably, our method excels in computational efficiency. Compared to MI and DMT requiring 208 and 410 seconds with $\times$23-45 overhead, our method runs in 70 seconds with minimal $\times$1.2 overhead to the base model (VC2).
Id. Leak $\downarrow$ Text $\uparrow$ MF $\uparrow$ Temp. C. $\uparrow$
DMT - 28.8 78.8 93.6
VMC - 27.1 93.7 94.6
MD - 30.4 93.9 93.3
MI - 31.1 95.5 93.5
MI 43.7 $\pm$ 1.5 29.2 $\pm$ 0.6 97.0 $\pm$ 0.4 96.8 $\pm$ 0.2
Ours 38.6 $\pm$ 1.8 28.8 $\pm$ 0.6 91.5 $\pm$ 0.9 97.0 $\pm$ 0.2
Table 1: Quantitative Evaluation Metrics, Motion Transfer. Values are reported as mean $\pm$ standard error of the mean (S.E.M).
Time Source Optim./ Infer Sum Over-
[sec.]: Invers. Tuning head $\downarrow$
Z.scope - - 9 9 -
DMT 260 opt&infer 150 410 $\times$45
MI 9 190 9 208 $\times$23
VC2 - - 58 58 -
Ours 12 0 58 70 $\times$1.2
Table 2: Runtime Comparison. Existing methods are $\times$23-45 slower than their base model. Ours adds merely $\times$1.2 overhead.
Qualitative Results: In Figure 4, we show that Motion by Queries helps guide generated videos to match the motion in source videos, including camera motion (top row), non-rigid object movement (middle row), and combinations of both types of motion (bottom row). While for quantitative results, we used Q injection ending at $t=600$. For qualitative results, we selected videos with Q injection ending between $t=800$ and $t=600$: closer to $t=800$ for camera motion and closer to $t=600$ for non-rigid movement. In Figure 5, we compare our results to MI and DMT. In the monkeys example, our method generates motion that resembles the source video, though it is with slightly lower motion magnitude compared to the baselines. For the boy and train examples, our results are comparable to those of MI and DMT. Additional comparisons reveal cases where the MI baseline suffers from identity leakage, aligning with the quantitative results. The supplemental also includes further comparisons and more examples of custom camera motion, and hybrid combinations of object and camera motion.
5 Application 2: Consistent multi-shot video generation
Generating videos from text prompts has progressed significantly, but current methods still struggle with maintaining long, coherent video sequences. A promising approach is to generate multiple short videos featuring the same characters, as cinematic videos often consist of shorter shots. The challenge lies in ensuring character consistency across scenes, as current T2V models excel at generating individual clips but struggle with consistency across multiple shots. Our goal is to generate multiple video shots with consistent characters, preserving their identity and motion dynamics across all scenes. This would open up new possibilities in content creation, storytelling, and entertainment.
Previous work  (Tewel et al., 2024; Fan et al., 2024) has focused on consistent character generation in the text-to-image domain by extending the self-attention mechanism, allowing images to share visual features. However, as shown in our analysis, video features often encode both identity and motion, leading to challenges when applying these methods to video. Naive extended attention mechanisms can cause synchronized or diminished motion, which makes image-based consistency methods unsuitable for video.
Building on this analysis, we propose a solution tailored for consistent video generation  [^1] The methods and results in this section are based on the arXiv version 1 (v1) Atzmon et al. (2024) of this work. Here, in version 2 (v2), we extend and further analyze those findings.. Our method preserves character identity and motion dynamics by combining extended attention for consistency with a vanilla, non-consistent text-to-video step to maintain dynamic motion. In the following, we describe the key modifications we introduced to Tewel et al. (2024) for consistent video generation.
ConsiStory Tewel et al. (2024) (See extended overview in the appendix, Sec. B.2) applies extended attention by using a subject mask, allowing images to share K, V self-attention features to maintain subject consistency. This reduces layout variability in the generated images. Then, to restore layout diversity, ConsiStory injects Q features from a vanilla, non-consistent sampling step. However, while this method improves diversity in images, it does not work well for videos, where the injected Q features fail to maintain the desired consistency and diversity across frames.
Refer to caption     Refer to caption
Figure 6: (click-to-view-online) Comparing Q token intervention strategies for consistent video generation. “Ours” (top row) balances character consistency and natural motion. VideoCrafter2 (second row) offers diverse motion but doesn’t allow for character consistency. “Full Q Preservation” (third row) without flow-based processing, preserving original motion but losing character consistency since identity leaks from VideoCrafter2. “No Q Intervention” (bottom row) maintains character consistency but suffers from motion degradation and synchronization across shots. The right side of each example shows y-t slices (temporal cross-sections) along the yellow vertical line visible in each frame, revealing motion patterns over time.
The naive extension of the extended attention approach to videos involves an extended self-attention mechanism that allows frames from different videos to attend to each other. However, this leads to reduced motion within each video. We hypothesize that self attention primarily ensures coherence within a video-shot, so extending it to multiple video-shots causes the different video-shots to become overall aligned, reducing the motion. Since the video Q-features are strongly tied to layout and motion, one approach to preserve each video’s original motion (i.e., pre-extended attention) is to inject Q-features from its vanilla version. While this helps preserve motion diversity, it also results in identity leakage from the original video, breaking our goal of maintaining character consistency across videos, as shown in Fig. 3.
To address this, we propose a two-stage solution for motion injection. Q preservation followed by Q Flow. In the early diffusion steps, which primarily control motion and layout, we inject the Q-features from vanilla videos generated without extended attention.
Then in the subsequent Q-Flow phase we apply a relaxed version of our Q-injection that preserves the flow of Q-features rather than injecting the features directly. This technique is inspired by TokenFlow Geyer et al. (2023), originally designed for T2I models.
Specifically, for each vanilla video, we first compute the nearest-neighbor correspondence field on its Q-features, defining a Q-feature flow that we aim to maintain in the generated consistent video. We then inject this correspondence into our generated video. For the exact definition of our Q Flow injection, see Appendix B.12.
5.1 Experiments
To demonstrate the importance of the steps suggested above, we conducted a study to investigate the effects of self-attention query (Q) tokens on motion and identity for consistent video generation. In the Appendix (Sec. B), we further show extensive evaluation including comparisons to baselines, a user study, and an ablation study. Fig. 6 illustrates typical generations for different interventions on Q tokens when combined with the extended self-attention mechanism of Fig. B.3. When we do not intervene in the Q tokens (Fig. 6, 4th row - “No Q Intervention”), subject identity is well-maintained across video shots, but motion quality significantly degrades. This manifests in: 1) Motion synchronization: movements become synchronized across video shots, e.g., the dog’s head turning simultaneously in all shots. 2) Reduced variability in motion style and pose: similar actions are repeated across shots, e.g., the dog’s leap, the Muppet’s centered swaying, the camera movement becomes static in the skating Muppet shot. 3) Motion artifacts: to reconcile the reduced motion variability with each scene’s text prompt, videos tend towards motion-artifacts. For example, the skating Muppet’s body appears frozen while its legs are displaced to visually accommodate the “skating” action. In contrast, combining extended attention with injected Q tokens cached from vanilla diffusion-sampled video shots (Fig. 6, 3rd row - “Full Q Preservation”) restores motion but largely loses subject identity. For instance, the Muppet’s colors revert to those of the vanilla model.
These observations reiterate our main finding about the dual nature of Q tokens. Injecting vanilla Q tokens restores motion, showing their influence on movement. Simultaneously, it leads to a leakage of the vanilla identity into the multi-shot consistency mechanism. This causes a loss of the shared subject identity (e.g., the Muppet’s color change).
Our approach (Fig. 6, 1st row - “Ours”) achieves a balance between both worlds. By intervening in the Q tokens throughout generation, we restore most of the original motion, including nuanced details like body and face orientations, postures, and natural movement of specific body parts (dog’s ears, Muppet’s hands and legs). Even the parallax style of video shots is preserved (right Muppet video shot). Our method’s effectiveness stems from our two-step process. First, Q preservation in early denoising steps establishes motion structure before identity is fully set. Then, flow-based Q injection allows Q values to evolve to better match the novel (more consistent) generation while enforcing some structural alignment through the flow process.
6 Limitations
In motion transfer, our method achieves slightly average motion fidelity compared to existing approaches, particularly with rapid movements (Figure 7 Appendix, top row). When transferring motion between different object classes, the source object’s shape can occasionally influence the target, such as when a pirate ship takes on characteristics of the source swan, or when a Colosseum inherits spatial properties from a mug (Figure 7).
In multi-shot consistent generation, balancing identity preservation and motion quality is still challenging. We find that Q injection may be too strong and still hurt identity. To manage this, motion preservation can be compromised by partially dropping out Q injection (See Section B.14).
7 Conclusion
Our analysis reveals key differences in how Q-injection behaves in text-to-video versus text-to-image models. Videos require more denoising steps to establish motion patterns, and Q vectors encode both motion and identity information, extending beyond their traditionally understood role in image models. When using extended attention for multi-shot video generation, motion tends to synchronize across videos, requiring longer Q-injection periods that affect both motion and identity. These insights enabled two practical applications: an efficient zero-shot motion transfer method and a training-free approach for consistent multi-shot videos. In a broader perspective, these findings challenge the separation of ”where” and ”what” pathways established in text-to-image models Tewel et al. (2023); Patashnik et al. (2023); Alaluf et al. (2023), providing novel evidence that queries in video models influence both motion (structure over time) and identity, suggesting a more complex interplay between these pathways than previously appreciated.
7.1 Broader Impact
This work advances text-to-video generation, offering applications in creative and entertainment industries. However, it introduces ethical challenges, including the potential misuse for creating misleading or harmful content. Future societal impacts could involve both enhanced accessibility to video creation and increased risks of misinformation.
Appendix A Appendix
Figure 7: Limitations. Cases where the source subject shape affects the target object.
A.1 Background: Self-Attention in T2V models
Appendix B Consistent Video Generation - Supplementary Detalis
B.1 Notations